OpenStudy (anonymous):

Solve lim n-> (infinity) = 1/n^3 n(n +1)(2n + 1)/(6)

OpenStudy (anonymous):

recheck the numerator, there is a gap between the first fraction.

OpenStudy (anonymous):

it looks like this hold on \[\lim_{n \rightarrow infinity} = 1/n^3 = n(n + 1)(2n + 1)/6 \]

OpenStudy (anonymous):

Do you mean this? \[\lim_{n \rightarrow \infty} {1 \over n^3}\cdot{n(n+1)(2n + 1) \over 6}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\lim_{n\rightarrow \infty} \frac{n(n+1)(2n+1)}{6n^3}\] \[\frac{1}{6} \lim_{n\rightarrow \infty} \frac{(n + 1)(2n+1)}{n^2}\] \[\frac{1}{6} \lim_{n \rightarrow \infty} \frac{2n^2 + 3n + 1}{n^2}\] \[\frac{1}{6} \lim_{n \rightarrow \infty} \left(2 + \frac{3}{n} + \frac{1}{n^2}\right)\] Which is \(\large {1 \over 6} * 2 = {1 \over 3}\)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i'll try the second one of my self thanks!

OpenStudy (anonymous):

Alright, if you get stuck you post it and I'll try to explain it :p

OpenStudy (anonymous):

ok so for this one i got lim n -> infinity = 1/n^2 n(n + 1)/2 is a function with a numerator and denominator degree 2 to the limit is ration of the leading coefficients, namely 1/2

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

@Meepi so i got that one right ?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

okay so the first one its 1/3 and this one 1/2

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

no problem :D